a solution has an [OH-] = 1.8x10^-10 at 25°C. what is the [H3O+] of the solution?
Answer #1

kw = [ka] [kg] kwa dierociation constant for water. 10-14 = [H3O+] [or .: [Hgot] = 10-14 - 1.8810-10 [Hgot] =5.55*10-5